Inventory Costing Methods—Perpetual Method

The Luann Company uses the perpetual inventory system. The following July data are for an item in Luann’s inventory:

July 1 Beginning inventory 30 Units@ $9 per unit
10 Purchased 50 Units@ $11 per unit
15 sold 60 Units
26 Purchased 25 Units@ $13 per unit

Calculate the cost of goods sold for the July 15 sale using (a) first-in, first-out, (b) last-in, first-out, and (c) the weighted-average cost methods.

Round your final answers to the nearest dollar. For weighted-average cost, do not round the weighted-average unit cost.

A. First-in, First-out:
cost of goods sold: $ ?
B. Last-in, first-out:
Cost of goods sold: $ ?
C. weighted-Avarge cost:
Cost of goods sold: $ ?
Answer #1

a) First in first out

Cost of goods sold = (30*9+30*11) = 600

b) Last in first out

Cost of goods sold = (50*11+10*9) = 640

c) Weighted average cost

Cost of goods sold = 820/80*60 = 615